Indicate the equation of the given line in standard form. The line with slope 9/7 and containing the midpoint of the segment whose end points are (2, -3) and (-6, 5).
Answer by sachi(548)

the midpoint of the segment whose end points are (2, -3) and (-6, 5)
is [(2-6)/2,(-3+5)/2]=(-2,1)
the equation of the line in standard form with slope 9/7 and containing (-2,1)
is (y-1)/(x+2)=9/7
or 7(y-1)=9(x+2)
or 7y-7=9x+18
or 7y-9x-25=0